# Yongyang Cai  Khyati Malik   Hyeseon Shin

### v2

@@ 6.3 Alternative Net Zero Scenarios
Along with the baseline emission cap scenario, we further analyze simulation results for the noncooperative model with alternative emission cap paths, defined by the net zero emission targets in 2050, 2070, and 2090. The top-left and top-right panels in Figure 8 display the emission permit prices and expected temperature increases under different emission cap scenarios. Under the net zero 2050 scenario, which is the most strict emission cap schedule, the emission permit price reaches \$1,593 per ton of carbon in 2049, and the temperature rise is restricted to 1.61 degrees Celsius by the end of this century. Net zero 2070 and net zero 2090 are more relaxed scenarios, leading to permit prices at \$445 and \$261 per ton of carbon in 2049, respectively. In these scenarios, the temperature rise by the end of the century is expected to reach 1.8 degrees Celsius and 1.99 degrees Celsius above the pre-industrial level, respectively. This result shows that the global target of restricting the temperature rise to 1.5 degrees Celsius is unattainable in a noncooperative world, even under the most restrictive net zero 2050 scenario, suggesting that stronger measures are needed to effectively regulate global emissions.
Figure 8: Simulation results of alternative emission cap scenarios under noncooperation with the ETS.
@@
The bottom-left and bottom-right panels in Figure 8 show the regional MAC and SCC, taking the US as an example. The comparison of the regional SCC and MAC for all other regions are available in Appendix A.8.5, which show the same patterns as the US. Our results show that stricter emission caps lead to higher MACs. Specifically, the MAC of the US under the net zero 2050 scenario can reach up to \$1,465 per ton of carbon in 2049, compared to the peak of \$1,323 in 2085 under the net zero 2090 scenario. This is because more rigorous emission caps imposed on each region entail additional abatement efforts, resulting in a higher MAC. We also find that more stringent emission caps result in a smaller SCC: the SCC of the US is \$218 per ton of carbon in 2100 in the net zero 2050 scenario, while it is \$198 in 2100 in the net zero 2090 scenario. This is because the permit price grows more quickly over time and even faster than the MAC (note that the optimal carbon tax is equal to the gap between the MAC and permit price when net emissions are strictly positive). Our finding suggests that under a stricter regulation on emission caps, carbon taxation becomes a less important policy instrument to control emissions.
@@ 7 Conclusion
In this work, we build a dynamic multi-region model of climate and the economy with a global emission cap-and-trade system. In our integrated assessment framework, 12 aggregated regions are allocated emission caps in line with the emission targets of the Paris Agreement and their net zero commitments. We solve for the market prices of emission permits under the dynamic Nash equilibrium in a noncooperative setting. The permit prices are endogenously determined by demand and supply of emission permits in the global permits market, reflecting regional heterogeneity in future productivity growth, abatement technologies, climate damage, and population growth. We further show that carbon taxation can be introduced under the global ETS as a complementary policy to help the social planner of each region achieve their optimal net emissions. For strictly positive net emissions, we have shown both theoretically and numerically that the optimal carbon tax is equal to the difference between the regional MAC and the market price of permits.
This work has several policy implications. First, our results indicate that the current global target of restricting the global temperature rise to 1.5 degrees Celsius above pre-industrial levels by 2100 is unattainable under noncooperation, even with the current emission commitments outlined in the Paris Agreement. Our findings suggest that more stringent emission reduction targets and global cooperation are needed to curb the trend of rising global temperature. Second, our baseline simulation shows that the current emission commitments of the Paris Agreement lead to excess emission permit supply in the initial decade, resulting in permit prices of zero. This finding suggests that effective implementation of the global ETS requires stricter emission caps so that the global supply of permits does not exceed the global demand of permits. Third, we demonstrate that the global ETS is not a pure substitute for the carbon tax, but instead that both can be implemented together to improve the efficiency of global climate policies.
@@ A.1 List of Parameters
Parameter Description
Economic parameters
$\beta=0.985$ Discount factor
$\gamma=1.45$ Elasticity of marginal utility
$\alpha=0.3$ Output elasticity of capital
$\delta=0.1$ Annual depreciation rate of capital
Climate parameters
$\zeta=0.0021$ Contribution rate of carbon emissions to temperature
$\sigma_{i,t}$ Carbon intensity (GtC per trillion \$)
$b_{1,i}$, $b_{2,i}$, $b_{3,i},$$b_{4,i}$ Abatement cost parameters
$\pi_{1,i}$, $\pi_{2,i}$ Climate damage parameters
TFP parameters
$g_{i,0},g_{i,t}$ Growth rate of TFP
$d_{i}$ Parameters of TFP growth
@@ A.2 List of Countries
Table A.2: List of countries for regional aggregation.
@@
Africa Algeria, Angola, Benin, Botswana, Burkina Faso, Burundi, Cameroon, Cape Verde, Central African Republic, Chad, Comoros, Democratic Republic of the Congo, Republic of the Congo, Cote d’Ivoire, Djibouti, Arab Republic of Egypt, Equatorial Guinea, Eritrea, Ethiopia, Gabon, Republic of The Gambia, Ghana, Guinea, Guinea-Bissau, Kenya, Lesotho, Liberia, Libya, Madagascar, Malawi, Mali, Mauritania, Mauritius, Morocco, Mozambique, Namibia, Niger, Nigeria, Rwanda, Sao Tome and Principe, Senegal, Seychelles, Sierra Leone, Somalia, South Africa, Sudan, Swaziland, Tanzania, Togo, Tunisia, Uganda, Zambia, and Zimbabwe.
European Union [^17] The current EU does not contain the United Kingdom, but in this paper we still assume the United Kingdom is in the EU for convenience. Austria, Belgium, Czech Republic, Denmark, Faeroe Islands, Finland, France, Germany, Greece, Greenland, Hungary, Iceland, Ireland, Italy, Luxembourg, Malta, Netherlands, Norway, Poland, Portugal, Slovak Republic, Spain, Sweden, Switzerland, Turkey, and the United Kingdom.
Eurasia Albania, Armenia, Azerbaijan, Belarus, Bosnia and Herzegovina, Bulgaria, Croatia, Estonia, Georgia, Kazakhstan, Kyrgyz Republic, Latvia, Lithuania, Macedonia FYR, Moldova, Montenegro, Romania, Serbia, Slovenia, Tajikistan, Turkmenistan, Ukraine, and Uzbekistan.
Latin America Antigua and Barbuda, Argentina, Aruba, Commonwealth of The Bahamas, Barbados, Belize, Bermuda, Bolivia, Brazil, Cayman Islands, Chile, Colombia, Costa Rica, Cuba, Dominica, Dominican Republic, Ecuador, El Salvador, Grenada, Guatemala, Guyana, Haiti, Honduras, Jamaica, Mexico, Netherlands Antilles, Nicaragua, Panama, Paraguay, Peru, Puerto Rico, St. Kitts and Nevis, St. Lucia, St. Vincent and the Grenadines, Suriname, Trinidad and Tobago, Uruguay, Bolivarian Republic of Venezuela, and the U.S. Virgin Islands.
Region Constituent Countries
Middle East Bahrain, Cyprus, Islamic Republic of Iran, Iraq, Israel, Jordan, Kuwait, Lebanon, Oman, Qatar, Saudi Arabia, Syrian Arab Republic, United Arab Emirates, West Bank and Gaza, and the Republic of Yemen.
Other Non-OECD Asia Afghanistan, Bangladesh, Bhutan, Brunei Darussalam, Cambodia, Fiji, French Polynesia, Indonesia, Kiribati, Democratic People’s Republic of Korea, Lao PDR, Malaysia, Maldives, Mongolia, Myanmar, Nepal, New Caledonia, Pakistan, Papua New Guinea, Philippines, Samoa, Solomon Islands, Sri Lanka, Thailand, Timor-Leste, Tonga, Vanuatu, and Vietnam.
Other High-Income Australia, Canada, Republic of Korea, New Zealand, Singapore.
A.3 Calibration of the TCRE Climate System
Each of the four RCP scenarios (Meinshausen et al., 2011) — RCP 2.6, RCP 4.5, RCP 6, and RCP 8.5 — provide their pathways of emissions, atmospheric carbon concentration, radiative forcing, and atmospheric temperature anomaly. When we calibrate the contribution rate of carbon emissions on temperature, $\zeta$, in a climate system, we use the pathways of emissions and atmospheric temperature anomaly of the four RCP scenarios. Figure A.1 shows that our calibrated TCRE climate system provides a very good projection of the atmospheric temperature anomaly (increase relative to pre-industrial levels) based on cumulative emissions only.
@@ A.3 Calibration of the TCRE Climate System
A.4 Emission Cap Pathways
Figure A.2 shows the regional emission cap pathways generated using the methodology described in Section 5.1. Figure A.3 compares the different emission cap pathways at the global level.
@@ A.4 Emission Cap Pathways
A.5 GDP Growth Rate beyond this Century

### v3

@@ 7.1 Alternative Emission Cap Scenarios
Along with the baseline emission cap scenario, we analyze simulation results for the noncooperative model with alternative emission cap paths, defined by the net zero emission targets in 2050, 2070, and 2090, where all regions achieve net zero emissions by the specified year. The top-left and top-right panels in Figure 6 display the emission permit prices and expected temperature increases under different emission cap scenarios. Under the net zero 2050 scenario, which is the most strict emission cap schedule, the emission permit price reaches \$1,539 per ton of carbon in 2049, and the temperature rise is restricted to 1.62°C by the end of this century. Net zero 2070 and net zero 2090 are more relaxed scenarios, leading to permit prices at \$514 and \$335 per ton of carbon in 2049, respectively. In these scenarios, the temperature rise by the end of the century is expected to reach 1.80°C and 1.98°C above the pre-industrial level, respectively. This result shows that the global target of restricting the temperature rise to 1.5°C is unattainable in a noncooperative world, even under the most restrictive net zero 2050 scenario, suggesting that stronger measures are needed to effectively regulate global emissions.
Figure 6: Sensitivity analysis: alternative emission cap scenarios.
@@
The bottom-left and bottom-right panels in Figure 6 show the regional MAC and SCC, taking the US as an example. The comparison of the regional SCC and MAC for all other regions are available in Appendix A.4.3, which show the same patterns as the US. Our results show that stricter emission caps lead to higher MACs. Specifically, the MAC of the US under the net zero 2050 scenario can reach up to \$1,471 per ton of carbon in 2048, compared to the peak of \$1,322 in 2086 under the net zero 2090 scenario. This is because more rigorous emission caps imposed on each region entail additional abatement efforts, resulting in a higher MAC. We also find that more stringent emission caps result in a smaller SCC: the SCC of the US is \$226 per ton of carbon in 2050 in the net zero 2050 scenario, while it is \$270 in 2050 in the net zero 2090 scenario. This is because the permit price in the net zero 2050 scenario grows more quickly over time and even faster than the MAC (recall that the optimal carbon tax is equal to the gap between the MAC and permit price when emissions are strictly positive). Our finding suggests that under a stricter regulation on emission caps, carbon taxation becomes a less important policy instrument to control emissions.
7.2 Alternative Climate Damages and Abatement Costs
Lastly, recognizing that climate damages and emission abatement costs are key drivers of our model outcomes, including permit prices, MAC, and the SCC, we conduct sensitivity analyses on the parameters for climate damages ($\pi_{1,i}$, $\pi_{2,i}$) and emission abatement ($b_{1,i}$, $b_{2,i}$, $b_{3,i},$ $b_{4,i}$). While we incorporate climate damage projections from Burke et al., (2018) as our baseline model simulation, we additionally consider projections from Kahn et al., (2021) and Nordhaus, 2010a . Specifically, we calibrate the climate damage parameters to match the projections from Kahn et al., (2021) and directly adopt the parameters from Nordhaus, 2010a . Figure 7 demonstrates the key economic and climate outcomes under different climate damage assumptions. The baseline climate damage parameters from Burke et al., (2018) result in slightly lower emission permit prices, reaching \$2,059 by 2069, compared to \$2,282 under the damage parameters estimated from Kahn et al., (2021) and \$2,294 under the damage parameters from Nordhaus, 2010a . With the same emission cap constraints imposed, the global temperature outcomes remain unaffected despite variations in the damage parameters. For the US, the baseline climate damage estimation from Burke et al., (2018) leads to higher MAC and higher SCC, indicating that the baseline marginal climate damages are relatively higher than those projected by Kahn et al., (2021) and Nordhaus, 2010a . However, regional heterogeneity exists; for example, Russia experiences negative SCC under the baseline parameter values, while experiencing small but positive SCC under the parameter values estimated from Kahn et al., (2021) and Nordhaus, 2010a . See Appendix A.4.4 for a comparison across all 12 regions.
Figure 7: Sensitivity analysis: alternative climate damage parameters.
For our sensitivity analysis of the emission abatement cost parameters, we consider the parameter values from Nordhaus, 2010a in addition to the baseline parameter values calibrated from Ueckerdt et al., (2019), both of which share the same functional form of abatement cost. As shown in Figure 8, the permit price rises to \$1,767 by 2069 under the abatement cost estimate of Nordhaus, 2010a , approximately 85% of the permit price projected under the baseline scenario. Despite the lower permit price, the global temperature increase remains similar to the baseline simulation, reaching 1.70°C by the end of the century, due to the emission cap constraints. The lower emission permit price under the abatement cost estimate of Nordhaus, 2010a reflects lower MAC, as illustrated with the US case in Figure 8, with similar patterns observed across most regions (see Appendix A.4.5). Lastly, the SCC is also lower under Nordhaus, 2010a parameters, with the SCC of the US reaching \$101 by 2100—just 32% of the baseline scenario.
Figure 8: Sensitivity analysis: alternative abatement cost parameters
@@ 8 Conclusion
In this work, we build a dynamic multi-region model of climate and the economy with a global emission cap-and-trade system. In our integrated assessment framework, 12 aggregated regions are allocated emission caps in line with the emission targets of the NDCs and net zero commitments, as established under the Paris Agreement and the Glasgow Pact. We solve for the market prices of emission permits under the dynamic Nash equilibrium in a noncooperative setting. The permit prices are endogenously determined by demand and supply of emission permits in the global permits market, reflecting regional heterogeneity in future productivity growth, abatement technologies, climate damage, and population growth. We further show that carbon taxation needs to be introduced under the global ETS as a complementary policy to help the social planner of each region achieve their optimal emissions. For strictly positive emissions, we show both theoretically and numerically that the optimal carbon tax is equal to the difference between the regional MAC and the market price of permits.
This work has several policy implications. First, our results indicate that the current global target of restricting the global temperature rise to 1.5°C above pre-industrial level by 2100 is unattainable under noncooperation, with the current emission commitments outlined in the Paris Agreement and the Glasgow Pact. Our findings suggest that more stringent emission reduction targets and global cooperation are needed to curb the trend of rising global temperature. Second, our baseline simulation shows that the current emission commitments lead to excess emission permit supply in the initial years, resulting in permit prices of zero. This finding suggests that effective implementation of the global ETS requires stricter emission caps so that the global supply of permits does not exceed the global demand of permits. Third, we demonstrate that the global ETS is not a pure substitute for the carbon tax, but instead that both can be implemented together to improve the efficiency of global climate policies.
@@ A.1 List of Parameters
Parameter Value Description
1) Economic system parameters (from Nordhaus, (2017))
$\beta$ $0.985$ Annual discount factor
$\gamma$ $1.45$ Elasticity of marginal utility
$\alpha$ $0.3$ Output elasticity of capital
$\delta$ $0.1$ Annual depreciation rate of capital
2) Climate system parameters
$\zeta$ $0.0021$ Contribution rate of carbon emissions to temperature
Table A.2 lists the values of the baseline abatement cost parameters calibrated from Ueckerdt et al., (2019). The values of carbon intensity at annual time steps will be provided upon request.
Table A.2: Abatement cost parameters (baseline) calibrated from Ueckerdt et al., (2019)
US EU Japan Russia Eurasia China
$b_{1,i}$ 0.462 0.477 0.750 0.292 0.347 0.328
$b_{2,i}$ 2.859 2.670 2.011 2.499 3.243 2.822
$b_{3,i}$ 9.920 5.832 2.492 7.625 7.966 7.189
$b_{4,i}$ 0.182 0.114 0.2 0.2 0.168 0.168
India MidEast Africa LatAm OHI OthAs
$b_{1,i}$ 0.594 0.455 0.665 0.286 0.347 0.602
$b_{2,i}$ 2.802 2.574 3.636 3.828 3.243 3.995
$b_{3,i}$ 6.336 11.205 6.558 11.496 7.966 6.518
$b_{4,i}$ 0.2 0.2 0.2 0.2 0.168 0.163
Table A.3 lists the calibrated values of the climate damage parameters used in the baseline analysis.
Table A.3: Climate damage parameters (baseline) calibrated from Burke et al., (2018)
US EU Japan Russia Eurasia China
$\pi_{1,i}$ 0.0842 0.0489 0.0090 -0.4169 0.2678 0.0003
$\pi_{2,i}$ 0.0096 0.0011 0.0748 0.3094 0.0002 0.0008
India MidEast Africa LatAm OHI OthAs
$\pi_{1,i}$ 0.0017 0.3595 0.1886 0.1801 0.0123 0.2161
$\pi_{2,i}$ 0.3276 0.0088 0.0764 0.0030 0.0044 0.0224
Table A.4 lists the values of the TFP parameters calibrated from Burke et al., (2018).
Table A.4: TFP parameters calibrated from Burke et al., (2018)
US EU Japan Russia Eurasia China
$g_{i,0}$ 0.0033 0.0089 0.0085 0.0170 0.0094 0.0345
$d_{i}$ 0.0011 0.0010 0.0010 0.0154 0.0010 0.0308
India MidEast Africa LatAm OHI OthAs
$g_{i,0}$ 0.0332 0.0093 0.0218 0.0134 0.0076 0.0221
$d_{i}$ 0.0151 0.0010 0.0013 0.0010 0.0010 0.0062
@@ A.2 List of Countries
Table A.5: List of countries for regional aggregation
@@
Africa Algeria, Angola, Benin, Botswana, Burkina Faso, Burundi, Cameroon, Cape Verde, Central African Republic, Chad, Comoros, Democratic Republic of the Congo, Republic of the Congo, Cote d’Ivoire, Djibouti, Arab Republic of Egypt, Ethiopia, Gabon, Gambia, The Ghana, Guinea, Guinea-Bissau, Kenya, Lesotho, Libya, Madagascar, Mali, Mauritania, Mauritius, Morocco, Mozambique, Namibia, Niger, Rwanda, Senegal, Sierra Leone, South Africa, Sudan, Swaziland, Tanzania, Togo, Tunisia, Uganda, Zambia, Zimbabwe.
European Union [^23] The current EU does not contain the United Kingdom, but in this paper we still assume the United Kingdom is in the EU for convenience.. Austria, Belgium, Bulgaria, Croatia, Czech Republic, Denmark, Finland, France, Germany, Greece, Hungary, Ireland, Italy, Latvia, Lithuania, Luxembourg, Netherlands, Poland, Portugal, Romania, Slovak Republic, Spain, Sweden, United Kingdom.
Eurasia Albania, Armenia, Azerbaijan, Belarus, Bosnia and Herzegovina, Estonia, Georgia, Kazakhstan, Kyrgyz Republic, FYR Macedonia, Moldova, Serbia, Slovenia, Tajikistan, Turkey, Ukraine, Uzbekistan.
Latin America Argentina, Bahamas, The Belize, Bolivia, Brazil, Chile, Colombia, Costa Rica, Cuba, Dominican Republic, Ecuador, El Salvador, Guatemala, Haiti, Honduras, Mexico, Nicaragua, Panama, Paraguay, Peru, Puerto Rico, Trinidad and Tobago, Uruguay.
Middle East Cyprus, Islamic Republic of Iran, Iraq, Israel, Jordan, Lebanon, Oman, Qatar, Saudi Arabia, United Arab Emirates
Other Non-OECD Asia Afghanistan, Bangladesh, Bhutan, Brunei Darussalam, Cambodia, Fiji, Indonesia, Malaysia, Mongolia, Nepal, Pakistan, Philippines, Samoa, Solomon Islands, Sri Lanka, Thailand, Vanuatu, Vietnam.
Other High-Income Australia, Canada, Iceland, Republic of Korea, New Zealand, Norway, Switzerland.
A.3 Details about Calibration and Data
A.3.1 Calibration of the TCRE Climate System
Each of the four RCP scenarios (Meinshausen et al.,, 2011) — RCP 2.6, RCP 4.5, RCP 6, and RCP 8.5 — provide their pathways of emissions, atmospheric carbon concentration, radiative forcing, and atmospheric temperature anomaly. When we calibrate the contribution rate of carbon emissions on temperature, $\zeta$, in a climate system, we use the pathways of emissions and atmospheric temperature anomaly of the four RCP scenarios. Figure A.1 shows that our calibrated TCRE climate system provides a very good projection of the atmospheric temperature anomaly (increase relative to pre-industrial levels) based on cumulative emissions only.
@@ A.3.1 Calibration of the TCRE Climate System
A.3.2 Calibration of Total Factor Productivity and Climate Damage
Figures A.2 and A.3 show that with our calibrated TFP and climate damage coefficients, the GDP per capita $y_{i,t}^{\mathrm{NoCC}}$ or $y_{i,t}$ matches well with the projected data $y_{i,t}^{\mathrm{BDD,NoCC}}$ or $y_{i,t}^{\mathrm{BDD}}$ from Burke et al., (2018), respectively, for all regions.
Figure A.2: Fitting GDP per capita under no climate impact. Lines represent GDP per capita under no climate impact from Burke et al., (2018); marks represent fitted GDP per capita.
Figure A.3: Fitting GDP under climate impact. Lines represent the GDP per capita under the climate impact of RCP 4.5 from Burke et al., (2018); marks represent fitted values.
A.3.3 Calibration of Climate Damage from Kahn et al., (2021)
For sensitivity analysis on climate damage parameters in Section 7.2, we calibrate the climate damage parameters $\pi_{1,i}$ and $\pi_{2,i}$ by considering projections on GDP loss across different climate scenarios in Kahn et al., (2021), which shows the percentage loss in GDP per capita by 2030, 2050, and 2100 under the RCP 2.6 and RCP 8.5 scenarios for China, EU, India, Russia, and the US. We use their method and data to project the percentage loss in GDP per capita ($\Delta_{i,t}^{\mathrm{RCP26}}$ and $\Delta_{i,t}^{\mathrm{RCP85}}$) every year from 2020 to 2114 under the RCP 2.6 and RCP 8.5 scenarios for each of our 12 regions, employing the baseline setup in Kahn et al., (2021). Specifically, $\Delta_{i,t}^{\mathrm{RCP26}}=1-y_{i,t}^{\mathrm{RCP26}}/y_{i,t}^{\mathrm{base}}$ and $\Delta_{i,t}^{\mathrm{RCP85}}=1-y_{i,t}^{\mathrm{RCP85}}/y_{i,t}^{\mathrm{base}}$, where $y_{i,t}^{\mathrm{RCP26}}$, $y_{i,t}^{\mathrm{RCP85}}$, and $y_{i,t}^{\mathrm{base}}$ are GDP per capita under RCP 2.6, RCP 8.5, and the baseline scenario, respectively. Thus from equation (17) we obtain $(\pi_{1,i},\pi_{2,i})$ by solving the following minimization problem for each region $i$:
$\min_{\pi_{1,i},\pi_{2,i}}\ \sum_{t=0}^{94}\left(\frac{1+\pi_{1,i}T_{t}^{\mathrm{RCP85}}+\pi_{2,i}\left(T_{t}^{\mathrm{RCP85}}\right)^{2}}{1+\pi_{1,i}T_{t}^{\mathrm{RCP26}}+\pi_{2,i}\left(T_{t}^{\mathrm{RCP26}}\right)^{2}}-\frac{1-\Delta_{t,i}^{\mathrm{RCP26}}}{1-\Delta_{t,i}^{\mathrm{RCP85}}}\right)^{2}.$ (A.1)
Here $T_{t}^{\mathrm{RCP26}}$ and $T_{t}^{\mathrm{RCP85}}$ are the global average temperature anomalies at time $t$ (deviation from the pre-industrial temperature) under the RCP 2.6 and RCP 8.5 scenarios. Figure A.3 shows, with our calibrated climate damage coefficients, the ratios of GDP per capita between RCP 2.6 and RCP8.5 from our model, matches well with the ratios in Kahn et al., (2021).
Figure A.4: Fitting climate damage parameters. Lines represent the ratios of GDP per capita between RCP 2.6 and RCP8.5 from Kahn et al., (2021); marks represent fitted ratios.
A.3.4 Regional Emission Cap Pathways
Figure A.5 displays the regional emission cap pathways, measured in Gigatonne of Carbon (GtC), for the baseline emission cap scenario, generated using the methodology described in Section 5.1.
@@ A.3.4 Regional Emission Cap Pathways
Table A.6 lists the regional emission caps for every region in five-year time steps. The regional emission caps at annual time steps will be provided upon request.
Table A.6: Regional emission cap pathways under the baseline scenario (unit: GtC)
2020 2025 2030 2035 2040 2045 2050 2055 2060 2065 2070
US 1.603 1.330 0.812 0.609 0.406 0.203 0.000 0.000 0.000 0.000 0.000
EU 1.081 0.923 0.617 0.463 0.309 0.154 0.000 0.000 0.000 0.000 0.000
Japan 0.308 0.260 0.197 0.148 0.099 0.049 0.000 0.000 0.000 0.000 0.000
Russia 0.694 0.674 0.621 0.517 0.414 0.310 0.207 0.103 0.000 0.000 0.000
Eurasia 0.464 0.450 0.390 0.324 0.259 0.193 0.126 0.071 0.030 0.015 0.000
China 3.433 3.370 3.061 2.551 2.042 1.532 1.023 0.513 0.004 0.002 0.000
India 0.940 0.773 0.295 0.259 0.222 0.185 0.148 0.111 0.074 0.037 0.000
MidEast 0.638 0.638 0.530 0.446 0.363 0.280 0.196 0.140 0.083 0.041 0.000
Africa 0.743 0.712 0.621 0.526 0.434 0.342 0.251 0.188 0.125 0.063 0.000
LatAm 0.815 0.736 0.598 0.480 0.362 0.244 0.126 0.094 0.063 0.031 0.000
OHI 0.613 0.538 0.347 0.260 0.173 0.087 0.000 0.000 0.000 0.000 0.000
OthAs 0.924 0.882 0.630 0.535 0.440 0.345 0.249 0.172 0.094 0.040 0.000
Figure A.6 presents a comparison of global emission cap pathways under various scenarios: the baseline scenario, net-zero by 2050, net-zero by 2070, and net-zero by 2090. In the net-zero scenarios, it is assumed that all countries achieve net-zero emissions by the respective target years.
@@
A.3.5 GDP Growth Rate beyond this Century